OpenStudy (anonymous):

i simplify 17-(-8)-9 and i got 34 is 34 the correct answer for this promblem?

OpenStudy (solomonzelman):

no, wrong.

OpenStudy (solomonzelman):

17-(-8)=17+8 and then -9

OpenStudy (anonymous):

do i add

OpenStudy (solomonzelman):

when there are two minuses, yes, you add.

OpenStudy (anonymous):

i keep getting 34

OpenStudy (solomonzelman):

17+8=25 25-9 is 16. you are subtracting 9 not adding.

OpenStudy (anonymous):

ok i got 16 is that right

OpenStudy (solomonzelman):

Yup! Right!

OpenStudy (anonymous):

i simplify 7-3x-4+8 and i got 1 is that correct?

OpenStudy (solomonzelman):

hold on, you have -3x in it, so the answer has to involve x, right?

OpenStudy (solomonzelman):

7-3x-4+8 change the order -3x-4+8+7 add like terms -4+8+7=11 so -3x+11

OpenStudy (anonymous):

so it would be 3x-3x?

OpenStudy (solomonzelman):

no it would be -3x+11. add like terms and you will get that too.

OpenStudy (anonymous):

here is a attach ment

OpenStudy (anonymous):

the last question is what i need help on an the one above it

OpenStudy (anonymous):